# Transferring Features Across Language Models With Model Stitching

### v2

@@ 1 Introduction
A popular perspective to view the internal computation of LLMs is through the lens of features, or functions of the input that serve a downstream purpose like predicting the next token or composing to form model circuits [Olah et al., 2020, Huben et al., 2024]. The Linear Representation Hypothesis (LRH) posits that these features are represented as directions in a high dimensional space [Park et al., 2024, Elhage et al., 2022] and sparse dictionary learning methods [Olshausen and Field, 1996, Faruqui et al., 2015, Serre, 2006] such as Sparse Autoencoders (SAEs) have recently gained popularity for interpreting models through this lens [Huben et al., 2024, Bricken et al., 2023, Gao et al., 2025, Kissane et al., 2024]. A second hypothesis, Strong Model Universality [Li et al., 2015], predicts that good models learn the same data representations.
@@ 2 Language Model Stitching
Table 1: Model-layer pairs that we stitch between and the associated downstream next token CE losses on OpenWebText. All activations are taken before the layer’s computation. We primarily stitch from middle residual layers but for specific use cases (e.g., steering) we stitch between $\nicefrac{{3}}{{4}}$ of the depth. The losses are computed stitching from $A$ to $B$, $B$ to $A$, $A$ to $B$ back to $A$, and $B$ to $A$ back to $B$. We calculate relative losses to the model that comes last and color code with blue and purple for model $A$ and model $B$ respectively.
Model-Layer Pair $A$ $B$ $A\to B$ $B\to A$ $A\to B\to A$ $B\to A\to B$
$A$: pythia-70m-deduped.3
@@ 4 Downstream Applications
Figure 3: (a) Evaluations of transferred probes stitching from pythia-70m-deduped to pythia-160m-deduped averaged over $8$ binary classification datasets. If the probe is retrained (orange), we almost recover ground truth performance (blue) across all probing $k$s and most datasets. Even if probe is not retrained, in most datasets we are able to probe significantly better than random (green). When directly probing on the residual stream, we find that transferring a probe trained on 70m-deduped (dotted) reaches similar accuracy to a probe trained on 160m-deduped (dashed) without retraining. (b) Response language steering vectors are able to be transferred between gemma-2-2b.20 and gemma-2-9b.33. From left to right, we chart the % of responses in the target language for no steering, ground truth steering, and steering using a transferred vector averaged over all languages $L$ in EuroParl and prompts. (c) The relative transfer gap distribution is bimodal with concentrations at 00 and $1$, implying the transfer steering works well for some languages but poorly for others.
@@ 5 Feature Analysis
Figure 4: (a) An overview of the feature analysis pipeline for a simple example where $2$ augmentations are generated. Structural features activate on all prompts (intersection) whereas semantic features only activate on some but not all prompts. (b) The semantic/structural classification reveals a divergence in the attribution correlation transferability metric for non-dead features. We plot the densities of both categories separately and relative density above. Structural features transfer more consistently but semantic features are more polarized (dominate the upper and lower percentiles).
@@ Acknowledgments
We would like to thank Michael Lepori, Jake Russin, and other members of the LUNAR Lab at Brown University for feedback at various stages of this project. We are also grateful to Neel Nanda for his valuable input during the early stages. AC would also like to thank Thomas Chang, Patrick Peng, and the CSCI2222 course at Brown for fruitful discussions about this work. AS acknowledges the support of armasuisse Science and Technology through a CYD Doctoral Fellowship.
@@ Appendix A Stitch Training Details
The stitch training methodology is quite minimal. For the training dataset, we collect activations at the desired layers in both models over the first $180$k samples of OpenWebText with a context size of $512$ ($128$ for Gemma due to compute constraints) and evaluate over the next $1$k samples with the same context size. We mask out all special tokens. The stitches themselves are two separate Linear layers (with biases initialized to 00) that map between the dimensions of $A$ and $B$. We use the Adam optimizer with a learning rate of 1e-4 and clip gradient norms to $1.0$. We found minimal sensitivity to learning rate schedule, so we just use a cosine annealing decay, and found that $2$ epochs is sufficient for convergence (though even $1$ is probably enough).
@@ Appendix C Zero-Shot SAE Evaluations
[…]
@@ Appendix D SAE Training and FLOP Estimation Details
Table 3: FLOPs estimates for various important procedures with caching activations. All SAEs are trained with sparsity $k=64$ and width $32$k. A full run is $4$B tokens ($120$k iterations) for the SAEs and $200$M ($36$k iterations) tokens for the stitch. The Pythia-70m SAE is trained post layer $2$ and the Pythia-160m SAEs are trained post layer $3$.
@@ D.2 Scaling Law Fit
Model/[Scratch, Stitch] $A$ $\beta$
@@ F.1 Specific Language Results
Clip(Tranfer PerfGround Truth $\text{Clip}\left(\frac{\text{Tranfer Perf}}{\text{Ground Truth Perf}},0,1\right),$ (16)
and we define $/0$ 0 as 00 and $\infty$ or $\infty$- ∞ are clipped to $1$ and 00 respectively.
We can compute a language frequency approximation by sentence-tokenizing a corpus of text and using langdetect to classify the language of each sentence, counting the frequencies. In Figure 11, we plot the clipped relative transfer gap against the language log frequency (computed over $100$k samples of OpenWebText) and identify a positive correlation with frequent languages having better transfer performance.
@@ H.2 Example Ablations and Classifications
Table 8: Examples of structural vs. semantic features based on feature centric evaluations. Drawn from pythia-70m-deduped pre-layer $3$ SAEs. Tokens are highlighted based on activation strength.
@@ Appendix I Attention Patterns
All experiments were run on single Quadro $6000$ / RTX3090309030903090 (both have 24GB VRAM) configurations. We estimate total GPU hours to be around 10000 in total for the project and most components of the experiments run contiguously for at most $1$ day.

### v3

# Transferring Linear Features Across Language Models With Model Stitching

@@ 1 Introduction
A popular perspective to view the internal computation of LLMs is through the lens of features, or functions of the input that serve a downstream purpose like predicting the next token or composing to form model circuits [Olah et al., 2020, Huben et al., 2024]. The Linear Representation Hypothesis (LRH) posits that these features are represented as directions in a high dimensional space [Park et al., 2024, Elhage et al., 2022]. However, the features had been empirically observed to appear in superposition (activating with interference), making interpretation of residual stream activations difficult. Sparse dictionary learning methods [Olshausen and Field, 1996, Faruqui et al., 2015, Serre, 2006] such as Sparse Autoencoders (SAEs) have recently gained popularity for interpreting models through this lens by disentangling dense representations into sparsely activating feature sets [Huben et al., 2024, Bricken et al., 2023, Gao et al., 2025, Kissane et al., 2024]. A second hypothesis, Strong Model Universality [Li et al., 2015], predicts that good models learn the same data representations.
@@ 2 Language Model Stitching
Table 1: Model-layer pairs that we stitch between and the associated downstream next token CE losses on OpenWebText. All activations are taken before the layer’s computation. We primarily stitch from middle residual layers but for specific use cases (e.g., steering) we stitch between $\nicefrac{{3}}{{4}}$ of the depth. The losses are computed stitching from $A$ to $B$, $B$ to $A$, AA to $B$ back to $A$, and $B$ to AA back to $B$. We calculate relative losses to the model that comes last and color code with blue and purple for model $A$ and model $B$ respectively.
Model-Layer Pair AA $B$ $A\to B$ $B\to A$ $A\to B\to A$ $B\to A\to B$
AA: pythia-70m-deduped.3
@@ 4 Downstream Applications
Figure 3: (a) Evaluations of transferred probes stitching from pythia-70m-deduped to pythia-160m-deduped averaged over 88 binary classification datasets. If the probe is retrained (orange), we almost recover ground truth performance (blue) across all probing $k$s and most datasets. Even if probe is not retrained, in most datasets we are able to probe significantly better than random (green). When directly probing on the residual stream, we find that transferring a probe trained on 70m-deduped (dotted) reaches similar accuracy to a probe trained on 160m-deduped (dashed) without retraining. (b) Response language steering vectors are able to be transferred between gemma-2-2b.20 and gemma-2-9b.33. From left to right, we chart the % of responses in the target language for no steering, ground truth steering, and steering using a transferred vector averaged over all languages $L$ in EuroParl and prompts. (c) The relative transfer gap distribution is bimodal with concentrations at 0 and $1$, implying the transfer steering works well for some languages but poorly for others.
@@ 5 Feature Analysis
Figure 4: (a) An overview of the feature analysis pipeline for a simple example where 22 augmentations are generated. Structural features activate on all prompts (intersection) whereas semantic features only activate on some but not all prompts. (b) The semantic/structural classification reveals a divergence in the attribution correlation transferability metric for non-dead features. We plot the densities of both categories separately and relative density above. Structural features transfer more consistently but semantic features are more polarized (dominate the upper and lower percentiles).
@@ Acknowledgments
We would like to thank Michael Lepori, Jake Russin, and other members of the LUNAR Lab at Brown University for feedback at various stages of this project. We are also grateful to Neel Nanda for his valuable input during the early stages. AC would also like to thank Thomas Chang, Patrick Peng, and the CSCI2222 course at Brown for fruitful discussions about this work. AS acknowledges the support of armasuisse Science and Technology through a CYD Doctoral Fellowship. This project was supported in part by a Young Faculty Award from the Defense Advanced Research Projects Agency Grant #D24AP00261. Ellie Pavlick is a paid consultant for Google Deepmind. The content of this article does not necessarily reflect that of the US Government or of Google and no official endorsement of this work should be inferred.
@@ Appendix A Stitch Training Details
The stitch training methodology is quite minimal. For the training dataset, we collect activations at the desired layers in both models over the first 180180k samples of OpenWebText with a context size of $512$ ($128$ for Gemma due to compute constraints) and evaluate over the next 11k samples with the same context size. We mask out all special tokens. The stitches themselves are two separate Linear layers (with biases initialized to 0) that map between the dimensions of $A$ and $B$. We use the Adam optimizer with a learning rate of 1e-4 and clip gradient norms to $1.0$. We found minimal sensitivity to learning rate schedule, so we just use a cosine annealing decay, and found that $2$ epochs is sufficient for convergence (though even $1$ is probably enough).
@@ Appendix C Zero-Shot SAE Evaluations
As basic evaluations of how well the feature spaces match, we can compute some core metrics of our zero shot transferred SAEs and compare them against the original SAEs. We report $L_{0}$, Fraction of Unexplained Variance (FUV), Delta loss, and dead features $\%$ in Table 2. Importantly, we note that the transferred SAEs are not perfect - one explanation of why is that the weight matrices of the transferred SAE (Equation 6) are still rank $\leq d_{A}$.
@@ Appendix D SAE Training and FLOP Estimation Details
Table 3: FLOPs estimates for various important procedures with caching activations. All SAEs are trained with sparsity $k=64$ and width 3232k. A full run is 44B tokens (120120k iterations) for the SAEs and 200200M (3636k iterations) tokens for the stitch. The Pythia-70m SAE is trained post layer $2$ and the Pythia-160m SAEs are trained post layer $3$.
@@ D.2 Scaling Law Fit
Model/[Scratch, Stitch] AA $\beta$
@@ F.1 Specific Language Results
$\text{Clip}\left(\frac{\text{Tranfer Perf}}{\text{Ground Truth Perf}},0,1\right),$ (16)
and we define $0/0$ as 0 and $\infty$ or -$\infty$ are clipped to $1$ and 0 respectively.
We can compute a language frequency approximation by sentence-tokenizing a corpus of text and using langdetect to classify the language of each sentence, counting the frequencies. In Figure 11, we plot the clipped relative transfer gap against the language log frequency (computed over 100100k samples of OpenWebText) and identify a positive correlation with frequent languages having better transfer performance.
@@ H.2 Example Ablations and Classifications
Table 8: Examples of structural vs. semantic features based on feature centric evaluations. Drawn from pythia-70m-deduped pre-layer 33 SAEs. Tokens are highlighted based on activation strength.
@@ Appendix I Attention Patterns
All experiments were run on single Quadro $6000$ / RTX30903090 (both have 24GB VRAM) configurations. We estimate total GPU hours to be around 10000 in total for the project and most components of the experiments run contiguously for at most $1$ day.